Compute the electric field experienced by a negative test charge given the following conditions in a vacuum: a. Source charge: –4.33 × 10^–12 C Distance from source charge: 2.33 × 10^–5 m (Your answer must be in this format : 1.11x10^1 N/C, -1.11x10^-1 N/C)


1
Expert's answer
2022-03-08T08:55:40-0500

The electric field

"E=k\frac{q}{r^2}\\
=9*10^9*\frac{4.33*10^{–12}}{(2.33*10^{-5})^2}=1.13*10^{12}\:\rm N/C"